The cost of 5 boxes of envelopes and 5 boxes of note paper is $ 25.95. Two boxes of envelopes and 6 boxes of note paper cost $ 22.38. Find the cost of each box of envelopes and each box of note paper.
Answer by ewatrrr(24785)


Hi,
let x represent the cost a box of envelopes and y the cost of a box of note paper.
5x + 5y = 25.95
5(x+y) = 25.95
x + y = 5.19
y= (5.19-x)
write as we read***
2x + 6(5.19-x)= 22.38
2x -6x + 31.14 = 22.38
31.14 - 22.38 = 4x
8.76 = 4x
$2.19 = x cost of a box of envelopes
(5.19 - 2.19 ) = $3.00 the cost of a box of note paper.
checking our answer
$4.38 + 6*$3.00 = $4.38 + $18.00 = $22.38